OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 Which of the following is a correct next step to solve the equation?

OpenStudy (texaschic101):

it would either be : x - 5 = 3 and solve for x or - (x - 5) = 3 and solve for x

OpenStudy (anonymous):

witch 1

OpenStudy (texaschic101):

either one....do you have answer choices ?

OpenStudy (texaschic101):

I would probably do the x - 5 = 3 first...but I don't think it matters...because there is gonna be 2 answers and that is how you find them

OpenStudy (anonymous):

ight